OpenStudy (anonymous):

WHAT ARE THE COORDINATES OF THE X INTERCEPT THAT HAVE THE POINTS (0,32) (100,212)

OpenStudy (amistre64):

i see this as the Celsius to Farenheit line ... hmmm

OpenStudy (amistre64):

the slope is ..... 100,212 -0, -32 -------- 100,180 = 180/100 = 9/5 y = \(9\over5\)x +32

OpenStudy (amistre64):

y = 0 when x = ... -32 = 9/5 x x = -32(5/9) = -160/9 ?

OpenStudy (anonymous):

WHICH ONE IS THE ANSWER

OpenStudy (amistre64):

thats right... (-160/9 , 0)

OpenStudy (amistre64):

do you have any choices to compare this to?

OpenStudy (anonymous):

NO

OpenStudy (amistre64):

do you know if you need a raounded decimal? or an exact answer?

OpenStudy (amistre64):

...rounded decimal that is....

OpenStudy (anonymous):

I DON'T THINK SO BECAUSE IT DOESN'T SAY TO ROUNDED OR DECIMAL

OpenStudy (amistre64):

then the exact qnswer would be:\[(-\frac{160}{9},0)\]

OpenStudy (amistre64):

there should be a fraction line bar beteen the 160 and the 9; but sometimes it gets lost ....

OpenStudy (anonymous):

THANK YOU ,AND I HAVE ANOTHER QUESTION FOR AN INPUT VALUE (X VALUE) OF 10, WHAT IS THE OUTPUT (Y) VALUE

OpenStudy (amistre64):

im assuming we are working with the line that is defined between these 2 points; y = (9/5)x + 32 ; when x = 10 y = 9(10/5) + 32 y = 9(2) + 32 y = 18 + 32 y = 50 ; when x = 10

OpenStudy (anonymous):

THANKS

OpenStudy (anonymous):

WHAT DOES REAL LIFE RELATIONSHIP DOES THE GRAPH REPRESENT